Please solve the simultaneous equation:
x/2+y/5=4
x/4+y/2=6





Answer by ewatrrr(24785)

HI,
*Note: First, multiply all terms of the second equation by -2, in order to eliminate one of the terms when the equations are added together.
.
-2*(x/4)+ -2*(y/2)=-2*6
.
Now, adding together with the first
(-x/2)+ -y = -12
x/2 + y/5 = 4
gives
(y/5) - y = -8
.
(-4/5)y = - 8
.
y = -8*(-5/4)
y = 10
.
x/2 + (10/5) = 4
x/2 + 2 = 4
x/2 = 2
x = 4